# Ira: Efficient Transaction Replay for Distributed Systems

### v1

@@ Abstract
We evaluated Ira-L against the state-of-the-art Ethereum client reth over two weeks of Ethereum mainnet activity ($100,800$ blocks containing over $24$ million transactions). Our hints are compact, adding a median of $47$ KB compressed per block ($\sim 5\%$ of block payload). We observe that the sequential hint generation and block execution imposes a $28.6\%$ wall-time overhead on the primary, though the direct cost from hints is $10.9\%$ of execution time; all of which can be pipelined and parallelized in production deployments. On the backup side, we observe that Ira-L achieves a median per-block speedup of $25\times$ over baseline reth. With $16$ prefetch threads, aggregate replay time drops from $6.5$ hours to $16$ minutes ($23.6\times$ wall-time speedup).
@@ 1 Introduction
The primary-backup paradigm forms the cornerstone of fault-tolerant distributed systems, from traditional databases [huntZooKeeperWaitfreeCoordination2010, junqueiraZabHighperformanceBroadcast2011] to modern blockchain networks [woodEthereumSecureDecentralised2014, yinHotStuffBFTConsensus2019]. In the standard State Machine Replication (SMR) model [schneiderImplementingFaulttolerantServices1990], a designated primary executes batches of transactions and propagates them to backup replicas, who must replay these transactions to maintain consistent state.
@@
Figure 1: LRU vs. Belady-optimal caching (cache size is 22). For access sequence $C$, AA, $C$ starting with cache $\{A,B\}$: LRU evicts AA (least recently used) when $C$ arrives, then misses when it accesses AA next. Optimal evicts $B$ (never accessed again), achieving one fewer miss. With hints, backups can make optimal eviction decisions.
@@ 1.1 Our Approach: Hint-Based Replay
Figure 2: Ira architecture. The primary executes transactions while building a hint that records access information. Backups use these hints to implement optimal cache eviction during replay.
In the case where the exact access pattern is known, the optimal solution is the Belady [beladyStudyReplacementAlgorithms1966, mattsonEvaluationTechniquesStorage1970] algorithm for caching, which informally states that the optimal caching algorithm is to evict the keys that are used furthest from the current operation. Implementing it is traditionally impossible as the access pattern can never be determined a priori.
Our key insight is that the inefficiency from Fig.˜1 stems from information asymmetry: the primary possesses complete knowledge of the access pattern but discards this information when communicating with backups. The primary can encode this knowledge compactly and transmit it as hints that enable backups to implement near-optimal replay. We propose Ira, a framework to accelerate replay through hint-based replay. Ira operates two components:
@@ 1.2 Contributions
    Ira for Ethereum. We present Ira-L, which achieves optimal cache performance through per-block hints that enable Belady-optimal eviction.
@@
Paper Organization. Section˜2 presents our system model and the two-pass protocol. Section˜4 describes the Ira-L algorithm. Section˜5 presents experimental results. Section˜6 discusses limitations and future work. Section˜7 surveys related work, and Section˜8 concludes.
@@ 2 Model and Preliminaries
Belady’s MIN Algorithm. Belady’s MIN algorithm [beladyStudyReplacementAlgorithms1966], proved optimal by Mattson et al. [mattsonEvaluationTechniquesStorage1970], provides the theoretically optimal cache replacement policy: when eviction is necessary, evict the item whose next access is furthest in the future. This policy minimizes total cache misses but requires knowledge of future accesses; information that is typically unavailable. In primary-backup replication, however, the primary has already executed the workload and possesses this knowledge. Ira exploits this observation: the primary encodes future access information as hints that enable backups to implement near-optimal cache replacement.
@@ 2.2 Ethereum: A Primer
Ethereum [woodEthereumSecureDecentralised2014] is a decentralized blockchain platform that executes arbitrary programs called smart contracts. We provide background on the aspects relevant to understanding Ira-L; Appendix˜A describes the execution client architecture in detail.
@@
Ethereum Virtual Machine (EVM). The EVM is a stack-based virtual machine that executes contract bytecode. State-accessing opcodes include:
@@ 2.3 Ira: Hint-Based Replay Architecture
In the standard primary-backup model, the primary executes transactions and observes the complete access pattern, for instance which keys are read and written, in what order, but discards this information when propagating transactions to backups. Backups must independently reconstruct the access pattern during replay, making cache eviction decisions without knowledge of future accesses. This information asymmetry leads to suboptimal performance: backups incur cache misses that the primary could have predicted and prevented.
@@
    Hint Construction: After execution, the primary constructs a hint $H$ containing the access set $A$, optionally annotated with metadata (e.g., where to read each key, access ordering for Belady-optimal eviction).
@@
  • Access ordering: The sequence in which keys are accessed, enabling Belady-optimal cache management when the access set exceeds cache capacity.
@@
The choice of hint granularity depends on the system’s I/O characteristics, network bandwidth, and memory constraints. Section˜4 presents Ira-L, a concrete instantiation that uses per-block access sets with source annotations for Ethereum block replay. Appendix˜D discusses how to adapt Ira to other primary-backup systems including Raft-based databases, streaming replication, and distributed key-value stores.
Fault Tolerance. This architecture is agnostic to the underlying consensus or replication protocol. We rely on the underlying protocol, whether Paxos [lamportPaxosMadeSimple2001], Raft [ongaroSearchUnderstandableConsensus2014], or Byzantine fault-tolerant protocols like PBFT [castroPracticalByzantineFault1999], for correctness guarantees. Ira is purely an optimization layer: hints do not affect the semantics of transaction execution or the safety properties of the replication protocol. A backup that receives incorrect or malicious hints may experience degraded performance (additional cache misses) but will never compute incorrect state.
@@ 3.1 Methodology
Dataset. We collected traces from blocks $24,019,447$ through $24,120,246$ ($100,800$ consecutive blocks spanning two weeks of Ethereum mainnet activity). The trace contains 1,192,888,6391,192,888,639{} state operations across approximately $15$ million transactions. We store traces in Apache Parquet [apachesoftwarefoundationApacheParquet] columnar format with Zstandard [facebookZstandardRealtimeData] compression, yielding a ${\sim}12$ GB dataset.
@@ 3.2 Findings
#2. Storage dominates I/O. Table˜2 and Figure˜4 show the distribution of state operations within the I/O operations in our trace. Storage operations (SLOAD and SSTORE) account for $75.6{}\%$ of all state accesses. Call operations, while numerous, result in negligible bytecode I/O ($\sim 1$ KB/block) because bytecode is immutable and cached effectively by the client. Account metadata queries comprise less than $1\%$ of operations.
@@
To isolate intra-block locality, we examine the distribution of per-key access counts within blocks. Figure˜5 shows this distribution: $34.5{}\%$ of keys are accessed only once within their block, while $65.5{}\%$ are accessed multiple times. The distribution exhibits a long tail, with some keys accessed up to 23,40623,406{} times within a single block.
@@
#4. Extreme Access Concentration. Access patterns exhibit extreme concentration at both the contract and key levels. Figure˜6 shows the cumulative storage access share as a function of contract rank: the top $3$ contracts (USDT, USDC, and Compound) account for $42.2{}\%$ of all storage operations, and the top $10$ contracts account for $63.6{}\%$. Table˜6 in the appendix provides the detailed breakdown.
@@
At the key level, concentration is even more pronounced. The top $1.23{}\%$ of keys (430,873430,873{} keys with $100+$ accesses) account for $84.0{}\%$ of all storage operations. Conversely, $64{}\%$ of keys are accessed only once or twice, contributing just $3.6{}\%$ of total accesses.
#5. Ephemeral Key Dominance. Despite high intra-block reuse, most keys are ephemeral and appear in only one block. Figure˜7 shows the distribution of key lifespans: $72.4{}\%$ of keys appear in only a single block, and $21.4{}\%$ appear in 2255 blocks. Only $1.0{}\%$ of keys persist across more than $50$ blocks. Table˜7 in the appendix provides the detailed breakdown.
@@
Consecutive block overlap, i.e., the fraction of keys in block $n+1$ that also appeared in block $n$, averages only $12.4{}\%$, has a median of $11.7{}\%$, and a max of $100.0{}\%$. This low overlap implies that standard cache replacement policies like LRU cannot effectively predict which keys will be needed in upcoming blocks.
#6. Bounded Working Set. Figure˜8 summarizes per-block statistics. The median block performs 8,6298,629{} total operations, of which 6,5876,587{} are storage operations accessing 1,8061,806{} unique keys. Even at the $95$th percentile, blocks access only 3,9203,920{} unique keys. Table˜8 in the appendix provides the detailed breakdown.
@@ 4 Ira-L: Latency-Optimal Algorithm
The workload analysis in Section˜3 reveals four properties that jointly motivate Ira-L’s design. I/O accounts for $67.9{}\%$ of block execution time (Finding #1), and storage operations dominate that I/O at $75.6{}\%$ of all state accesses (Finding #2). The per-block working set is bounded, with a median of 1,8061,806{} unique keys (Finding #6), making per-block hints tractable. Yet $72.4{}\%$ of keys are ephemeral, appearing in only a single block (Finding #5), and consecutive-block overlap averages just $12.4{}\%$. Hence, LRU and other heuristic replacement policies cannot effectively predict which keys the next block will need. The primary, however, already knows the complete access set for each block it executes. Discarding this information creates an unnecessary information asymmetry between primary and backup.
Ira-L eliminates this asymmetry. After executing block $b$, the primary constructs a hint $H_{b}$ containing three components: storage keys with source annotations, account addresses, and bytecode addresses. The primary transmits $H_{b}$ alongside the block data. The backup uses $H_{b}$ to prefetch all required state before execution begins, eliminating cold misses entirely.
@@ 4.1 Hint Generation
Without source annotations, the backup must perform a two-step lookup for every key: first seeking the history index (a B-tree traversal on a 60-byte composite key, followed by roaring bitmap decompression), then reading the value from the appropriate table. The source annotation eliminates the first step for Plain State keys, reducing two seeks to one and eliminating all I/O for Zero keys. Across our dataset, $46.6{}\%$ of storage keys resolve to PlainState, $10.5{}\%$ to Zero, and $43.0{}\%$ to Changeset (Section˜C.1). In total, $57.0{}\%$ of keys bypass the history index entirely, with a per-block median of $55.3{}\%$. The source hint removes the confirmation step for PlainState keys and all I/O for Zero keys; only Changeset keys require the full two-step lookup. Appendix˜A describes the full history-index lookup algorithm.
@@ 4.2 Hint-Based Replay
This pipelining ensures that the backup’s wall-clock time is bounded by $\max(\text{prefetch},\text{execution})$ rather than their sum. Appendix˜A provides implementation details of the threading model and channel sizing.
@@ 4.3 Correctness and Safety
Advisory Hints. Hints affect only performance, not correctness. If hints are missing, corrupted, or stale, the backup can fall back to standard execution without prefetching, incurring the same I/O cost as the baseline, but producing correct state. This property enables incremental deployment: nodes that do not support hints operate normally, while hint-aware nodes benefit from faster replay. Incorrect hints may cause degraded performance (cache misses trigger the fallback path) but never produce incorrect state.
@@ 5.1 Configuration
Methodology. We evaluate Ira-L using the same two-week dataset of 100,800100,800{} consecutive Ethereum mainnet blocks (blocks 24,019,44724{,}019{,}44724,120,24624{,}120{,}246). All benchmarks start with a cold OS page cache (purged before each run) to measure worst-case I/O performance. We compare four configurations:
@@
  • Backup (parallel-$k$): backup replaying with $k$ concurrent prefetch threads ($k\in\{16,64\}$).
@@ 5.2 Hint Overhead on Primary
The primary generates hints during normal block execution. We measure two components of overhead: hint construction (building the next-access map) and hint serialization (compressing and writing hints to disk). We compare against the unmodified baseline to quantify the total cost.
Wall-Time Overhead. Table˜3 summarizes the end-to-end wall-time comparison. The baseline executes all 100,800100,800{} blocks in 23,25023,250{} s, while the primary completes in 29,88929,889{} s which is a $28.6{}\%$ increase in wall time. This overhead reflects not only hint generation but also any secondary effects of instrumentation on execution (e.g., additional memory pressure, cache interference).
Table 3: Wall-time comparison between baseline and primary.
Configuration Wall Time (s) Overhead
Baseline 23,25023,250{} -
Primary (with hints) 29,88929,889{} $28.6{}\%$
Per-Block Hint Cost. To isolate the direct cost of hint generation, we examine the per-block breakdown of hint construction and serialization time as a fraction of execution time (Figure˜9). Across all blocks, hint construction consumes 940940{} s and hint serialization 778778{} s, totaling 1,7191,719{} s of hint-related work, which is $10.9{}\%$ of the aggregate execution time 15,71715,717{} s. The median per-block hint fraction is $12.1{}\%$, with a $95$th-percentile value of $22.1{}\%$.
Figure 9: CDF of per-block hint cost as a percentage of block execution time. Construction and serialization each contribute roughly half the total overhead, with a median of $12.1{}\%$ and a $95$th percentile of $22.1{}\%$.
@@
[…]
These per-block fractions (median $12.1{}\%$, P95 $22.1{}\%$) reflect a conservative, fully sequential measurement where hint construction and serialization run inline with block execution. In practice, both components are independent of execution correctness and can be overlapped with the next block’s processing. Hint construction ($6.0{}\%$ of aggregate execution time) builds a map from keys already collected during EVM execution and queries the storage history index to annotate each key with its optimal read source (Appendix˜A). Hint serialization ($5.0{}\%$) compresses and persists the hint to disk, requiring no executor state at all. Production reth already employs a multi-threaded architecture with a dedicated executor thread pool and a separate storage writer thread (Appendix˜A); in this model, hint construction can execute on the existing thread pool and serialization can run on the storage thread, overlapping both with the next block’s execution. Under such pipelining, the effective overhead approaches zero for all but the final block in a batch. Moreover, the cost is paid once by the primary and recovered on every backup: even with sequential prefetching, the backup replays $5.2{}\times$ faster in wall time (Section˜5.3), yielding a net system-wide speedup despite the primary’s $28.6{}\%$ wall-time increase.
Hint Sizes. Figure˜11 shows the distribution of per-block hint sizes. The median compressed hint is 46.546.5{} KB (mean 51.551.5{} KB), with a $95$th-percentile of 104.2104.2{} KB and a maximum of 239.0239.0{} KB. Zstandard compression achieves a $2.17{}\times$ reduction from the raw representation (median 102.2102.2{} KB raw). Figure˜12 confirms that these sizes remain stable across the entire 100,800100,800{}-block range, with no visible trend or drift.

### v2

@@ Abstract
We evaluated Ira-L against the state-of-the-art Ethereum client reth over two weeks of Ethereum mainnet activity ($100,800$ blocks containing over $24$ million transactions). Hint generation adds $10.9\%$ overhead to primary execution time. On the backup, hint-driven prefetching speeds up aggregate replay by $5.2\times$ with a single prefetch thread and by $23.6\times$ with $16$ threads. Our hints are compact, adding a median of $47$ KB compressed per block (${\sim}5\%$ of block payload).
@@ 1 Introduction
The primary-backup paradigm forms the cornerstone of fault-tolerant distributed systems [grayDangersReplicationSolution1996], from traditional databases [huntZooKeeperWaitfreeCoordination2010, junqueiraZabHighperformanceBroadcast2011] to modern blockchain networks [woodEthereumSecureDecentralised2014, yinHotStuffBFTConsensus2019]. In the standard State Machine Replication (SMR) model [schneiderImplementingFaulttolerantServices1990], a designated primary executes batches of transactions and propagates them to backup replicas, who must replay these transactions to maintain consistent state.
@@
Figure 1: LRU vs. optimal caching (cache size is 22). For access sequence $C$, AA, $C$ starting with cache $\{A,B\}$: LRU evicts AA (least recently used) when $C$ arrives, then misses when it accesses AA next. Optimal evicts $B$ (never accessed again), achieving one fewer miss. With hints revealing the access set, backups can prefetch all required keys before execution, eliminating misses entirely.
@@ 1.1 Our Approach: Hint-Based Replay
Figure 2: Ira architecture. The primary executes transactions while building a hint that records the access set. Backups use these hints to prefetch all required state before replay, achieving zero cache misses.
When the cache is smaller than the working set, Belady’s MIN algorithm [beladyStudyReplacementAlgorithms1966, mattsonEvaluationTechniquesStorage1970] provides the optimal eviction policy: evict the key whose next access is furthest in the future. However, when the cache can hold the entire working set, a simpler and more effective strategy is prefetching: load all required keys before execution begins, eliminating misses entirely. In primary-backup replication, the primary already knows the complete access set—this knowledge enables prefetching on backups.
Our key insight is that the inefficiency from Figure˜1 stems from information asymmetry: the primary possesses complete knowledge of the access pattern but discards this information when communicating with backups. The primary can encode this knowledge compactly and transmit it as hints that enable backups to implement near-optimal replay. We propose Ira, a framework to accelerate replay through hint-based replay. Figure˜2 illustrates the architecture. Ira operates two components:
@@ 1.2 Contributions
    Ira for Ethereum. We present Ira-L, which achieves zero-miss cache performance through per-block hints that enable complete working-set prefetching.
@@ 2 Model and Preliminaries
Cache Optimization with Future Knowledge. Belady’s MIN algorithm [beladyStudyReplacementAlgorithms1966], proved optimal by Mattson et al. [mattsonEvaluationTechniquesStorage1970], provides the theoretically optimal cache replacement policy: when eviction is necessary, evict the item whose next access is furthest in the future. However, when the cache can hold the entire working set, a simpler strategy—prefetching—is more effective: load all required keys before execution begins, eliminating misses entirely. Both strategies require knowledge of future accesses, information that is typically unavailable. In primary-backup replication, however, the primary has already executed the workload and possesses this knowledge. Ira exploits this observation: the primary encodes access information as hints that enable backups to prefetch state or implement optimal eviction, depending on the relationship between working set size and cache capacity.
@@ 2.2 Ethereum: A Primer
Ethereum [woodEthereumSecureDecentralised2014] is a decentralized blockchain platform that executes arbitrary programs called smart contracts. The execution client runs the Ethereum Virtual Machine (EVM), a stack-based virtual machine that interprets contract bytecode. We provide background on the aspects relevant to understanding Ira-L; Appendix˜C describes the execution client architecture in detail.
@@
The EVM is a stack-based virtual machine that executes contract bytecode. State-accessing opcodes include:
@@ 2.3 Ira: Hint-Based Replay Architecture
In the standard primary-backup model, the primary executes transactions and observes the complete access pattern—which keys are read and written, in what order—but discards this information when propagating transactions to backups. Backups must independently discover the access pattern during replay, making cache eviction decisions without knowledge of future accesses. This knowledge gap leads to suboptimal performance: the primary possesses future access information that backups lack, causing backups to incur cache misses that the primary could have predicted and prevented.
@@
    Hint Construction: After execution, the primary constructs a hint $H$ containing the access set $A$, optionally annotated with metadata (e.g., where to read each key, or access ordering if cache capacity is limited).
@@
  • Access ordering: The sequence in which keys are accessed, enabling optimal cache eviction when the access set exceeds cache capacity.
@@
The choice of hint granularity depends on the system’s I/O characteristics, network bandwidth, and memory constraints. Section˜4 presents Ira-L, a concrete instantiation that uses per-block access sets with source annotations for Ethereum block replay. Appendix˜G discusses how to adapt Ira to other primary-backup systems including Raft-based databases, streaming replication, and distributed key-value stores.
Fault Tolerance. This architecture is agnostic to the underlying consensus or replication protocol. We rely on the underlying protocol, whether Paxos [lamportPaxosMadeSimple2001], Raft [ongaroSearchUnderstandableConsensus2014], or Byzantine fault-tolerant protocols like PBFT [castro1999practical], for correctness guarantees. Ira is purely an optimization layer: hints do not affect the semantics of transaction execution or the safety properties of the replication protocol. A backup that receives incorrect or malicious hints may experience degraded performance (additional cache misses) but will never compute incorrect state. Appendix˜B discusses incentive mechanisms for encouraging correct hint generation.
@@ 3.1 Methodology
Dataset. We collected traces from blocks $24,019,447$ through $24,120,246$ ($100,800$ consecutive blocks spanning December 15–29, 2025 on Ethereum mainnet). The trace contains 1,192,888,6391,192,888,639{} state operations across approximately $15$ million transactions. We store traces in Apache Parquet [apachesoftwarefoundationApacheParquet] columnar format with Zstandard [facebookZstandardRealtimeData] compression, yielding a ${\sim}12$ GB dataset.
@@ 3.2 Findings
#2. Storage dominates I/O. Table˜2 shows the distribution of state operations in our trace (Figure˜17 in the appendix shows the per-block breakdown). Storage operations (SLOAD and SSTORE) account for $75.6{}\%$ of all state accesses. Call operations, while numerous, contribute minimal I/O overhead: contract bytecode is loaded once per contract and cached in memory, so repeated calls to the same contract incur no disk I/O. Account metadata queries comprise less than $1\%$ of operations.
@@
To isolate intra-block locality, we examine the distribution of per-key access counts within blocks. Figure˜4 shows this distribution: $34.5{}\%$ of keys are accessed only once within their block, while $65.5{}\%$ are accessed multiple times. The distribution exhibits a long tail, with some keys accessed up to 23,40623,406{} times within a single block.
@@
#4. Extreme Access Concentration. Access patterns exhibit extreme concentration at both the contract and key levels. Figure˜5 shows the cumulative storage access share as a function of contract rank: the top $3$ contracts (USDT, USDC, and Compound) account for $42.2{}\%$ of all storage operations, and the top $10$ contracts account for $63.6{}\%$. Table˜6 in the appendix provides the detailed breakdown.
@@
At the key level, concentration is even more pronounced. The top $1.23{}\%$ of keys (430,873430,873{} keys with $100+$ accesses) account for $84.0{}\%$ of all storage operations. Conversely, $64{}\%$ of keys are accessed only once or twice, contributing just $3.6{}\%$ of total accesses. This concentration might suggest that caching the most frequently accessed keys globally would suffice. However, as Finding #5 shows, $72.4{}\%$ of keys are ephemeral (appearing in only one block), and consecutive-block overlap averages just $12.4{}\%$. The “hot” keys change unpredictably from block to block, making static caching ineffective.
#5. Ephemeral Key Dominance. Despite high intra-block reuse, most keys are ephemeral and appear in only one block. Figure˜6 shows the distribution of key lifespans: $72.4{}\%$ of keys appear in only a single block, and $21.4{}\%$ appear in 2255 blocks. Only $1.0{}\%$ of keys persist across more than $50$ blocks. Table˜7 in the appendix provides the detailed breakdown.
@@
Consecutive block overlap, i.e., the fraction of keys in block $n+1$ that also appeared in block $n$, averages only $12.4{}\%$, has a median of $11.7{}\%$, and a max of $100.0{}\%$. This low overlap implies that standard cache replacement policies like LRU cannot effectively predict which keys will be needed in upcoming blocks: LRU retains recently-used keys, but with only $12.4{}\%$ overlap, most cached keys will not be accessed in the next block, while most keys needed will not be in cache.
#6. Bounded Working Set. Figure˜7 summarizes per-block statistics. The median block performs 8,6298,629{} total operations, of which 6,5876,587{} are storage operations accessing 1,8061,806{} unique keys. Even at the $95$th percentile, blocks access only 3,9203,920{} unique keys. Table˜8 in the appendix provides the detailed breakdown.
@@
Takeaway. Our analysis reveals a fundamental tension. On one hand, I/O dominates execution ($67.9{}\%$), storage dominates I/O ($75.6{}\%$), and per-block working sets are bounded (median 1,8061,806{} keys), making caching critical and per-block hints tractable. On the other hand, $72.4{}\%$ of keys are ephemeral and consecutive-block overlap averages just $12.4{}\%$, so LRU and other heuristic policies cannot predict which keys the next block will need. The key insight is that this unpredictability is artificial: the primary already knows the complete access set for each block it executes, but discards this information. This creates an unnecessary information asymmetry between primary and backup.
@@ 4 Ira-L: Latency-Optimal Algorithm
Ira-L eliminates the information asymmetry identified in Section˜3 by having the primary share its access knowledge with backups. After executing block $b$, the primary constructs a hint $H_{b}$ containing three components: storage keys with source annotations, account addresses, and bytecode addresses (defined in Section˜4.1). The primary transmits $H_{b}$ alongside the block data. The backup uses $H_{b}$ to prefetch all required state before execution begins, eliminating cold misses entirely.
@@ 4.1 Hint Generation
Without source annotations, the backup must perform a two-step lookup for every key: first seeking the history index (a B-tree traversal on a 60-byte composite key, followed by roaring bitmap decompression), then reading the value from the appropriate table. The source annotation eliminates the first step for Plain State keys, reducing two seeks to one and eliminating all I/O for Zero keys. Across our dataset, $46.6{}\%$ of storage keys resolve to PlainState, $10.5{}\%$ to Zero, and $43.0{}\%$ to Changeset (Section˜E.2). In total, $57.0{}\%$ of keys bypass the history index entirely, with a per-block median of $55.3{}\%$. The source hint removes the confirmation step for PlainState keys and all I/O for Zero keys; only Changeset keys require the full two-step lookup. Appendix˜C describes the full history-index lookup algorithm.
@@ 4.2 Hint-Based Replay
This pipelining ensures that the backup’s wall-clock time is bounded by $\max(\text{prefetch},\text{execution})$ rather than their sum. Appendix˜C provides implementation details of the threading model and channel sizing.
@@ 4.3 Correctness and Safety
Advisory Hints. Hints affect only performance, not correctness. If hints are missing, corrupted, or stale, the backup can fall back to standard execution without prefetching, incurring the same I/O cost as the baseline, but producing correct state. This property enables incremental deployment: nodes that do not support hints operate normally, while hint-aware nodes benefit from faster replay. Incorrect hints may cause degraded performance (cache misses trigger the fallback path) but never produce incorrect state. Detection and mitigation of corrupted hints is discussed in Appendix˜B.
@@ 5.1 Configuration
Methodology. We evaluate Ira-L using the same two-week dataset of 100,800100,800{} consecutive Ethereum mainnet blocks (blocks 24,019,44724{,}019{,}44724,120,24624{,}120{,}246). All benchmarks start with a cold OS page cache (purged before each run) to measure worst-case I/O performance. We compare four configuration families:
@@
  • Backup (parallel-$k$): backup replaying with $k$ concurrent prefetch threads ($k\in\{1,2,4,6,8,10,12,16,32,64\}$). We report detailed per-block results for $k=16$ and $k=64$; the full sweep is used in the thread-scaling analysis (Section˜5.4).
@@ 5.2 Hint Overhead on Primary
The primary generates hints during normal block execution. We measure two components of overhead: hint construction (collecting the access set) and hint serialization (compressing and writing hints to disk). We compare against the unmodified baseline to quantify the total cost.
Wall-Time Overhead. Table˜3 summarizes the overhead. The baseline executes all 100,800100,800{} blocks in 23,25023,250{} s, while the primary completes in 29,88929,889{} s—a $28.6{}\%$ increase. Of the hint-related work, construction (collecting the access set and querying the history index) accounts for 940940{} s ($6.0{}\%$ of aggregate execution time) and serialization (compressing and writing to disk) accounts for 778778{} s ($5.0{}\%$), totaling 1,7191,719{} s ($10.9{}\%$). The remaining wall-time increase reflects secondary effects of instrumentation (e.g., memory pressure, cache interference).
Table 3: Primary overhead breakdown. Wall time compares end-to-end execution; hint cost decomposes the overhead into construction and serialization as a fraction of aggregate execution time (15,71715,717{} s).
Component Time (s) % of Exec Time
Baseline wall time 23,25023,250{}
Primary wall time 29,88929,889{}
     Hint construction 940940{} $6.0{}\%$
     Hint serialization 778778{} $5.0{}\%$
     Hint total 1,7191,719{} $10.9{}\%$
Wall-time overhead $28.6{}\%$
Per-Block Hint Cost. Figure˜8 shows the per-block distribution of hint cost as a fraction of execution time. The median per-block hint fraction is $12.1{}\%$, with a $95$th-percentile of $22.1{}\%$. Although construction accounts for more aggregate time than serialization (Table˜3), the serialization CDF has a heavier right tail: serialization time is roughly constant (${\sim}12.5{}$ ms), so it dominates the per-block fraction for short-executing blocks, while construction scales with block size and its fraction remains bounded. Each CDF is sorted independently, so the total-cost curve is the distribution of per-block sums, not the pointwise sum of the two component curves.
Figure 8: CDF of per-block hint cost as a percentage of block execution time. Each curve is an independent distribution: the total-cost curve is the CDF of per-block $(t_{\text{construct}}+t_{\text{serialize}})/t_{\text{exec}}$, not the sum of the two component CDFs. Median total hint fraction is $12.1{}\%$; P95 is $22.1{}\%$.
@@
These per-block fractions (median $12.1{}\%$, P95 $22.1{}\%$) reflect a conservative, fully sequential measurement where hint construction and serialization run inline with block execution. In practice, both components are independent of execution correctness and can be overlapped with the next block’s processing. Hint construction ($6.0{}\%$ of aggregate execution time; Table˜3) builds a map from keys already collected during EVM execution and queries the storage history index to annotate each key with its optimal read source (Appendix˜C). Hint serialization ($5.0{}\%$) compresses and persists the hint to disk, requiring no executor state at all. Production reth already employs a multi-threaded architecture with a dedicated executor thread pool and a separate storage writer thread (Appendix˜C); in this model, hint construction can execute on the existing thread pool and serialization can run on the storage thread, overlapping both with the next block’s execution. Under such pipelining, the effective overhead approaches zero for all but the final block in a batch. Moreover, the cost is paid once by the primary and recovered on every backup: even with sequential prefetching, the backup replays $5.2{}\times$ faster in wall time (Section˜5.3), yielding a net system-wide speedup despite the primary’s $28.6{}\%$ wall-time increase.
Hint Sizes. Figure˜9 shows the distribution of per-block hint sizes. The median compressed hint is 46.546.5{} KB (mean 51.551.5{} KB), with a $95$th-percentile of 104.2104.2{} KB and a maximum of 239.0239.0{} KB. Zstandard compression achieves a $2.17{}\times$ reduction from the raw representation (median 102.2102.2{} KB raw). Figure˜10 confirms that these sizes remain stable across the entire 100,800100,800{}-block range, with no visible trend or drift.